Hank purchased a $25,500 car two years ago using a 7.2 percent, 3-year loan. He has decided that he would sell the car now, if he could get a price that would pay off the balance of his loan.
What’s the minimum price Hank would need to receive for his car? (Round the loan payment to the nearest cent, but do not round any other interim calculations. Round your final answer to 2 decimal places.)

the correct answer is not 9,116.94. so please help me out
Answer #1

PV of Loan =25500
Number of Years =3
The annual payment =PV/((1-(1+r)^-n)/r) =25500/((1-(1+7.2%)^-3)/7.2%) =9752.3438

The balance of amount after 2 years =PV*(1+r)^n-PMT*((1+r)^n-1)/r) =25500*(1+7.2%)^2-9752.3438*((1+7.2%)^2-1)/7.2%) =
9097.34